Al is seven years older than Maria. The sum of their ages is 51. What is each of their ages?
Answer by stanbon(75887)

Al is seven years older than Maria. The sum of their ages is 51. What is each of their ages?
-------
Equations:
a = m + 7
a + m = 51
------
Substitute for "a" and solve for "m"::
m+7 + m = 51
2m = 44
m = 22 (maria's age now)
-----
Solve for "a"::
a = m + 7
a = 22+7
a = 29 (Al's age now)
-----------
Cheers,
Stan H.
------------